OpenStudy (anonymous):

MEDALS FOR WHO SOLVES THE EQUATIONS! Substitution by elimination •7. 2x – 3y = 4 and 3x + 2y = 6 Answer: (2,0) •8. 2x – 5y = 6 and 4x + 3y = -1 Answer: (1/2,1) How do you solve to get these answers. Show work

OpenStudy (jdoe0001):

have you cover doing system of equations yet?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 2x - 3y = 4&{\color{red}{ \times 2}}\implies &4x\cancel{-6y}=8\\ 3x + 2y = 6&{\color{red}{ \times 3}}\implies &9x\cancel{+6y}=18\\ \hline\\ &&\square x\qquad =\square \end{array}\) that'd give you "x", then substitute on either equation to get "y"

OpenStudy (anonymous):

do you add or subtract.

OpenStudy (jdoe0001):

you add, then again, if the number is negative, the adding will turn to subtraction anyway

OpenStudy (anonymous):

okay. So it would be 13x=26

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

which means x = 26/13

OpenStudy (anonymous):

x=2

OpenStudy (jdoe0001):

then you can grab that "x" and put it on either equation, solve for "y"

OpenStudy (jdoe0001):

2x-3y = 4 2(2)-3y=4 => 4-3y=4

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 2x - 5y = 6&{\color{red}{ \times -2}}\implies &-4x+10y=-12\\ 4x + 3y = -1&&\quad 4x + 3y = -1\\ \hline\\ &&\qquad \quad 13 y=-13 \end{array}\)

OpenStudy (jdoe0001):

and you do the same on that one, once one of the variables has been ELIMINATED, then the rest is just solving for the other

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 2x - 5y = 6&{\color{red}{ \times -2}}\implies &\cancel{-4x}+10y=-12\\ 4x + 3y = -1&&\quad \cancel{4x} + 3y = -1\\ \hline\\ &&\qquad \quad 13 y=-13 \end{array}\\\quad\\ {\color{blue}{ 13}}y=-13\implies y=\cfrac{-13}{{\color{blue}{ 13}}}\implies y=-1\)

OpenStudy (anonymous):

how do you get 1/2 for x?

OpenStudy (jdoe0001):

well we dunno what "x" is yet since we know "y", we'll use that in either equation, say the 2nd one so \(\bf 4x+3{\color{red}{ y}}=-1\implies 4x+3({\color{red}{ -1}})=-1\implies 4x-3=-1\)

OpenStudy (jdoe0001):

anyhow, that gives an "x", but is not 1/2 though

OpenStudy (anonymous):

okay thanks.

OpenStudy (jdoe0001):

yw